OpenStudy (anonymous):

find the equation of line L. L is parallel to y=3x. What is the equation of line L? points;(3,4) y=3x. can someone please help me solve this problem.

OpenStudy (anonymous):

y=3 x +b

OpenStudy (anonymous):

4 = 3(3) + b 4=9+b b=-5 y=3x -5